Choose a strategy to solve each problem.


Derek gave Kyle half of his marble collection. Kyle gave Janis half of the marble he received fom Derek. Janis kept 12 of the marbles and gave the remaining 16 to Jill. How many marbles did Derek give to Kyle?

Answer by ptaylor(2198)


Let x= Number of marbles in Derek's marble collection
Derek gave Kyle x/2 marbles
Kyle gave Janis (1/2)(x/2) or x/4 marbles
Now we are told in a roundabout way that Janis received 28 marbles, so:
x/4=28 multiply each side by 4
x=112------Number of marbles in Derek's marble collection

x/2=112/2=56------------------number of marbles that Derek gave Kyle
(1/2)*56=28-----Number of marbles that Kyle gave Janis and this checks.

Hope this helps---ptaylor